OpenStudy (anonymous):

Find the inverse of 2x - 7 and also, find f^-1(-2).. when it said f^-1(-2), i have to substitute -2 on the inverse of 2x -7 right?

jimthompson5910 (jim_thompson5910):

that is correct, so what is the inverse?

OpenStudy (anonymous):

y + 7 / 2 ?? :)

jimthompson5910 (jim_thompson5910):

remember you're swapping x and y, then solving for y so you should get (x+7)/2 which means it looks like you have it pretty much

jimthompson5910 (jim_thompson5910):

so f^(-1)(x) = (x+7)/2 f^(-1)(-2) = (-2+7)/2 ... plug in x = -2 f^(-1)(-2) = ???

OpenStudy (anonymous):

ahhh, okayyyy...5/2 :)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks! :)

jimthompson5910 (jim_thompson5910):

you're welcome